An alpha particle has a positive charge that is double the charge of a proton, and a mass of 6.64 ✕ 10−27 kg. At a particular instant, it is moving with a speed of 4.95 ✕ 106 m/s through a magnetic field. At this instant, its velocity makes an angle of 59° with the direction of the magnetic field at the particle's location. The magnitude of the field is 0.170 T.

(a) What is the magnitude of the magnetic force (in N) on the particle?

__________N

(b) What is the magnitude of the particle's acceleration (in m/s2) at this instant?

__________m/s2
Answer #1

ANSWER:::

a) Charge of alpha particle= q = double charge of proton

q= 2×1.6×10-19 C =3.2×10-19 C

  

Mass of particle = m= 6.64 ✕ 1027 kg.

Speed of v=4.95 ✕ 106 m/s   

and  =  angle of 59°

  The magnitude of the field is 0.170 T.= B

We know that,

F=q×V×B×sin

  

  

the magnitude of the magnetic force (in N) on the particle IS  ,

b) we know that

  

  

  

  

The magnitude of the particle's acceleration (in m/s2) at this instant is